11 May 2015
Last day of science – tomorrow we concentrate on packing
Logging of the deep NEEM hole was terminated last night at 4. The results are good, and Dorthe is eager to analyze them.
The GPS team went to a strain net point 25 km away and measured position.
Most of the cargo in camp is now collected in a row of stacks close to the old apron. The cargo is sorted according to what can be stowed and what we need to have access to on the traverse.
The mechanics garage is the last structure outside the dome, and it was emptied today. We plan to take it down tomorrow.
Two short cores were drilled through the roof of the balloon trench, and the show a hard, dense and uniform firn. Sepp has measured temperature in the 170 m borehole at the GPS main reference station.
What we have done today:
1. Visited GPS strain net point 25 km away.
2. Drilled cores through roof of balloon trench.
3. Packing down in mechanics garage.
4. NEEM deep borehole logging successfully completed.
5. Cleaning up in camp.
6. Repaired the winch on the GrIT Case tractor.
7. Organizing antennae and radios for the traverse.
8. Contact with our colleagues on the Renland Ice cap at 19 local (21 Renland
Ad.2: The cores show quite uniform firn with a density of 600 g/m3.
Ad.8: The Renland team reports all well. Weather is fine, and today they assembled the freezer for the ice cores. The drill tent is being outfitted with furniture and equipment, and the inclined trench for the drill was excavated. A melt layer (most likely from last year) was encountered at 1.7 m depth. Main drilling will probably begin the day after tomorrow. The shallow drill system has been setup and the team expect to begin the U.S. firn air drilling and gas collection tomorrow.
Weather: Fog in the morning, later clearing, low clouds in the evening. Temp. – 21 °C to -34 °C, 3-5 m/s from S. Visibility: In the morning less than 1 km, later unrestricted.
FL, J.P. Steffensen
|← Previous entry||Next entry →|